# RelianceScope: An Analytical Framework for Examining Students’ Reliance on Generative AI Chatbots in Problem Solving

### v1

@@ 7. Limitation and Future Work
We note several limitations. First, students’ reliance behaviors in our study may not fully generalize to interactions with commercial chatbots (e.g., ChatGPT), as our system employed customized prompts and students were aware that their interactions were logged. Although we designed appropriate incentives and clearly communicated learning objectives, time constraints and grading policies may have influenced students’ reliance behaviors. Second, we do not establish strong statistical evidence linking reliance patterns to learning gains. Noise in self-reported measures and the use of a one-time knowledge assessment limited our ability to fit robust regression models. Establishing such relationships will require larger datasets, repeated measures, and more diverse learner populations. We encourage future work to apply RelianceScope across varied and controlled learning settings to replicate and extend our findings and to more precisely characterize when and how reliance on AI supports learning.

### v2

@@ 7. Limitation and Future Work
We note several limitations. First, students’ reliance behaviors in our study may not fully generalize to interactions with commercial chatbots (e.g., ChatGPT), as our system employed customized prompts and students were aware that their interactions were logged. Although we designed appropriate incentives and clearly communicated learning objectives, time constraints and grading policies may have influenced students’ reliance behaviors. Second, we do not establish strong statistical evidence linking reliance patterns to learning gains. Noise in self-reported measures and the use of a one-time knowledge assessment limited our ability to fit robust regression models. Establishing such relationships will require larger datasets and repeated measures. We encourage future work to apply RelianceScope across varied and controlled learning settings to replicate and extend our findings and to more precisely characterize when and how reliance on AI supports learning.
Acknowledgements.
This work is funded by NSF Grant IIS-2442990. The findings and conclusions expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.